OpenStudy (anonymous):

I am asked to prove, using mathematical induction, that \[3 | (n^3 - n) for all positive integers. I am having a difficult time relating the basis step to the induction step when a operator other than equivalence is used. How would I approach this problem most effectively.

OpenStudy (anonymous):

Sorry my equation didn't' come out in the wash. \[3 | (n^3 -n )\]

OpenStudy (phi):

what's the basis step?

OpenStudy (mr.math):

Write \(n^3-n=3k\) for some integer \(k\). This would make it easier.

OpenStudy (anonymous):

@phi would assume the basis step is to determine that the formula is true for integer k. \[3 | (1^3 - 1) \\ 3 | ( 1 - 1 ) \\ 3 | 0\]

OpenStudy (phi):

now assume the statement is true for n. Show it is true for N+1 n^3-n = n(n^2-1)= (n-1)n(n+1) replace n with N+1 N(N+1)(N+2) N(N+1)( (N-1) + 3) can you finish?

OpenStudy (anonymous):

I see your logic @Mr.Math! I understand that finding equivlency is nesscary to proving the formula. I am just unsuring of my footing when making assumptions or introducing forumlas.

OpenStudy (mr.math):

Your fist step is right, you found that the statement is valid for \(n=1\), Now do as phi said. Assume it's true for \(n=m\), then show that this assumption implies that it's also true for \(n=m+1\).

OpenStudy (mr.math):

The factorization that phi did above is the key to the problem.

OpenStudy (anonymous):

I can easily grasp the induction step when I am trying to simplify a Left Hand Side to a Right Hand Side. This question has thrown me for a loop not having that option. I am unsure what I am supposed to make my induction step equivlent to.

OpenStudy (phi):

You have to show 3 | (n^3 - n) You first prove this is true for a base case, n=1 You then assume it is true for some n. Now show it is true for n+1. plug (n+1) into the formula and rewrite (see previous post) as n(n+1)( (n-1) + 3) = (n-1)n(n+1) + 3n(n+1) = (n^3-n) + 3n(n+1) 3 divides evenly into the first term by hypothesis. (We assumed the statement is true for n) The 2nd term is evenly divisible by 3 because it has 3 as a factor therefore the entire expression is divisible by 3. We conclude that if the statement is true for n, it will be true for n+1. But having proved it is true for n=1, we conclude it is true for n=2, and then n=3, and so on.

OpenStudy (anonymous):

Alright! I can see the connection now. All of my examples dealt with sumations where I could find an equivlence to prove ideas. Maybe this was a poor way of introducing the idea becuase I felt bound by this idea that I had to find equivlence when you provided a clear example that is any approptriate condition can used. I really appreciate your help and clarification.